Please explain me Principle of atom conservation in detail and the kind of questions that can be asked from this topic.

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear rishabh,

The principle implies Number of atoms remains constant. Lets understand this with an example.

1g of a mixture of cuprous oxide and cupric oxide was quantitatively reduced to 0.839 g of metallic copper. what is the weight of cupric oxide in the original sample.

The solution is:

63.563 + 16 = 79.563
127.126 + 16 = 143.126
x*79.563 + y*143.126 = 1
x*63.563 + y*127.126 = 0.839
x = 0.006925497
x*79.563 = 0.551 g cupric oxide

 

Please approve the answer if you like it.


Feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Rishabh Dabral
13 years ago

Sir,

I didn't understand the last three steps of the example....plz explain me in detail.